# Distribution Grids May Be a Barrier to Residential Electrification

### v2

@@
Replacing fossil-fueled appliances and vehicles with electric alternatives can reduce greenhouse gas emissions and air pollution in many settings. However, residential electrification can raise electricity demand beyond the safe limits of electrical infrastructure, increasing the risk of blackouts or requiring grid reinforcement that can be slow and expensive. Here, we estimate the physical and economic impacts on distribution grids of electrifying all housing and personal vehicles in each county of the lower 48 United States. We find that space heating is the main driver of grid impacts, with the coldest regions seeing demand peaks up to five times higher than today’s peaks. Accommodating electrification of all housing and personal vehicles could require up to 544 GW of distribution grid reinforcement nationally, at a cost of \$320 to \$720 billion, or \$2,600 to \$5,800 per household (95% confidence intervals). However, demand-side management can mitigate demand peaks, reducing grid reinforcement costs by up to 77%.
@@
Distribution grid, Electrification, Heat pumps, Electric vehicles
@@ INTRODUCTION
Fossil-fueled driving, space heating, and water heating contribute about one-third of United States greenhouse gas emissions 1. Replacing fossil-fueled vehicles and heaters with electric alternatives could significantly reduce these emissions today 2, and essentially eliminate them if the United States achieves its goal of 100% carbon-free electricity by 2035 3. For these reasons, the Inflation Reduction Act of 2022 incentivizes adoption of electric vehicles and heat pumps (efficient electric heating and cooling machines) through tax credits and point-of-sale rebates 4, 5, 6. State and local governments have enacted further incentives, including some bans on fossil-fueled vehicles or heaters 7. Due in part to these policies, adoption of electric vehicles and heat pumps in the United States is accelerating 8.
@@
While electrification of fossil-fueled activities could deeply reduce emissions, it could also have serious impacts on the power grid. Today, electricity demand in most United States power systems peaks in the hottest hours of the year, driven mainly by air conditioning 9. Electrifying space heating could heighten annual demand peaks and shift them from summer to winter 10, 11, 12, 13, 14. Electric vehicle adoption could further exacerbate demand peaks 15, 16, 17. Heightened demand peaks from electrification could require replacing power lines and transformers, which must be sized to accommodate worst-case power flows. Transformers in particular are already in short supply, with United States utilities experiencing lead times of up to two years and prices that have increased by 400 to 900% in the last three years 18. Utilities pass grid reinforcement costs on to ratepayers in the form of increased electricity prices 19, 20, which could weaken economic incentives for further electrification.
@@
Existing research on nationwide grid impacts of electrification focuses mainly on high-voltage transmission systems. The National Renewable Energy Laboratory’s Electrification Futures Study 21, 10, 22 investigates how electrification might alter load profiles and transmission infrastructure needs. In their high-electrification 2050 scenario, peak electricity demand rises 38% relative to a reference scenario. Wilson et al. 2 use the ResStock tool 23 to investigate cost and emission impacts under various heat pump adoption scenarios. While Wilson et al. 2 do not explicitly consider grid impacts, they stress the need for future work to understand how evolving load profiles should shape infrastructure investments.
@@
In parallel with nationwide high-voltage transmission grid research, several studies investigate medium-voltage distribution grid impacts in California specifically. Notably, Elmallah et al. 24 use the Distribution Deferral Opportunity Report, which the California Public Utilities Commission requires utilities to file each year, to analyze electrification impacts within the Pacific Gas & Electric service territory. Elmallah et al. estimate that distribution grid reinforcement to accommodate new load from electrification, including up to 100% adoption of electric vehicles and up to 50% adoption of heat pumps and water heaters, could cost \$1 to \$10 billion, or about \$200 to \$2,000 per customer. Li et al. 25 use the same dataset to study adoption of electric vehicles (but not heat pumps or water heaters) statewide, estimating that 67% of distribution feeders will need reinforcement, at a total cost of \$6 to \$20 billion or about \$300 to \$900 per customer. Several other state-specific studies have investigated the impacts of electrification on greenhouse gas emissions without considering impacts on power grids 14, 26, 27.
@@
As a complement to nationwide transmission-level studies and California-specific distribution-level studies, this paper investigates the impacts of electrification on distribution grids nationwide. In addition to expanding the scope of distribution-level studies from the 58 counties in California to 3,100 counties in the lower 48 United States, this paper investigates strategies to mitigate grid impacts through demand-side management of space heating, water heating, and personal vehicles. These are the three main drivers of grid impacts from electrification, and we are aware of no distribution-level study that has considered demand-side management strategies for all of them. We use the open-source EDGIE (Emulating the Distribution Grid Impacts of Electrification) toolbox 28 and make all data publicly available to facilitate reproduction or extension of the work. As illustrated in Fig. 1, EDGIE models heat pumps, electric vehicles, water heaters, and building thermal dynamics. It uses real data on weather 29, indoor temperature setpoints 30, domestic hot water use 31, and miscellaneous electrical loads 32. EDGIE implements empirically validated, multi-physics modeling at hourly time resolution and county-level spatial resolution. We tune building, appliance, and vehicle parameters to county-specific data. We find that almost all United States counties could require grid reinforcement in a 100% electrification scenario. In colder climate zones, peak demand rises by 200% on average relative to today’s peak demand, and up to 800% in the coldest counties under the worst-case equipment choices. Distribution grid reinforcement to accommodate 100% electrification nationwide could cost \$320 to \$720 billion, or about \$2,500 to \$5,800 per household. However, demand-side management strategies, such as improving equipment efficiencies and coordinating device operation, could reduce nationwide grid reinforcement costs by up to 77%.
@@ RESULTS
Figure 2: Peak aggregate electricity demand from 1,000 households in a hot week in Miami (left) barely exceeds today’s peak. In a cold week in Minneapolis (right), however, demand peaks at more than double today’s peak.
@@
We assess the physical impacts of home and vehicle electrification on distribution grids through bottom-up modeling. In each county, we simulate all the residential appliances and personal vehicles from 1,000 representative households, then rescale the aggregate load profile to reflect the true number of households. Simulated appliance and vehicle properties, housing sizes and types (detached single-family vs. attached multifamily), and levels of insulation and air sealing all vary both across and within counties. We simulate peak heating and cooling weeks under two scenarios: (1) A business-as-usual scenario that reflects today’s building stock, appliances, and vehicles, and (2) an all-electric scenario that reflects a possible future with complete electrification of personal vehicles and residential space and water heating. In the all-electric scenario, each home has a resistance water heater, an air-to-air heat pump for space heating and cooling, and backup resistance heat. We compute design heating and cooling loads for each home under EnergyStar heating and cooling design conditions 33. Each heat pump is sized to meet the larger of its design heating load and design cooling load, capped at the largest residential air-to-air heat pump size currently available. In the two northernmost climate zones, we simulate adoption of heat pumps that meet the United States Department of Energy’s Cold Climate Heat Pump Challenge specifications 34. We fit physical and behavioral parameters associated with buildings, thermal equipment, and vehicles to county- and climate-specific data, using real data wherever possible. The Experimental Procedures section and the Supplementary Information document provide more details on models and data sources.
@@
Figure 3: Today (top left), per-household demand peaks are typically higher in hotter areas and in cold rural areas with significant amounts of electric heating. In an all-electric future (top right), per-household peaks increase and shift to colder areas. Per-household grid reinforcement requirements (bottom) are highest in the Rocky Mountains, the upper Midwest, and the Northeast.
@@
The modeling approach outlined above enables analysis of the sizes and causes of likely grid impacts from electrification. Fig. 2 shows the aggregate power used by 1,000 households over a peak cooling week in Miami, Florida (left plot; a hot, humid climate) and a peak heating week in Minneapolis, Minnesota (right; cold). In these plots, the shaded areas from bottom to top represent miscellaneous electrical loads (blue; lights, computers, dishwashers, etc.), water heaters (orange), electric vehicles (yellow), and air-source heat pumps with backup resistance heat (purple). Dashed red lines show today’s annual demand peaks. In Miami, electrification roughly doubles peak demand relative to today’s peak. Electric vehicle adoption increases electricity demand, but the new air-to-air heat pumps – which can both heat and cool – offset some of this increase by improving air conditioning efficiency. In Minneapolis, by contrast, peak demand increases by about a factor of five, driven mainly by space heating in the coldest weather, when heat demand rises and heating efficiencies fall.
@@
We analyze grid impacts nationwide by running simulations for each county in the lower 48 United States. Fig. 3 shows heat maps of the estimated distribution grid capacity, normalized by the number of homes, in each county under the business-as-usual scenario (top left) and the all-electric scenario (top right). We estimate distribution grid capacities in each county and scenario by simulating the appliances and vehicles for 1,000 households, rescaling the aggregate demand to the true number of households, taking the 99th percentile, then adding 20% to reflect an infrastructure sizing safety margin. The top left plot shows that the normalized grid capacity (in units of kW per household) in the business-as-usual scenario tends to be higher in areas with more cooling demand, reaching a maximum of 13 kW per household in the hot, sunny Southwest. By contrast, the top right plot shows that the normalized grid capacity in the all-electric scenario is higher in colder areas, reaching a maximum of 25 kW per household in the northern Midwest.
@@
The lower heat map in Fig. 3 shows the estimated grid reinforcement capacity, normalized by the number of households, required to accommodate 100% electrification in each county. We estimate the grid reinforcement capacity requirement as the difference between the all-electric and business-as-usual grid capacities. As expected, northern regions require more grid reinforcement due to higher space heating demand. Grid reinforcement requirements reach a maximum of 24 kW per household in the northern Midwest. Many southern areas with mild winters require little or no grid reinforcement. In some cases, the lower heat map in Fig. 3 shows significant variation in grid reinforcement requirements between adjacent counties. This variation can be explained by differences in housing types and sizes, levels of insulation and air sealing, driving patterns, weather extremes, or other location-specific data. These geographic differences suggest a need for tailored infrastructure planning, as approaches that work in one county may not necessarily work in neighboring counties.
@@
To estimate the economic costs of reinforcing distribution grids to accommodate electrification, we multiply the aggregate grid reinforcement requirement for each county (in units of kW) by a \$/kW price. Informed by several recent studies, we choose this price to reflect a typical distribution grid upgrade. In reality, grid reinforcement prices vary significantly across the United States due to variation in grid topology and capacity, prevalence of overhead vs. underground power lines, costs of labor and equipment, and other factors. Due to a lack of reliable location-specific data on grid reinforcement prices, however, we use the same price distribution for each county.
@@
Figure 4: Statewide distribution grid reinforcement costs (top row) reflect both population densities and per-household grid reinforcement requirements. Per-household costs (bottom row) do not reflect population densities. Left column: Low-price scenario for grid reinforcement. Middle: Average price. Right: High price.
@@
We draw on four studies 24, 35, 36, 37 to determine a plausible range of prices for typical distribution grid reinforcement, which may include installing power lines, transformers, switchgear, capacity banks, or other infrastructure. Based on Pacific Gas & Electric data, Elmallah et al. 24 estimate prices of 860 to 36,500 \$/kW, with the highest prices for the smallest projects. (We inflation-adjusted all price data to 2024 dollars.) Larson et al. 35 use Energy Information Administration data 38 to derive state-wise prices and a national average of 1,664 \$/kW, including a variety of administrative costs. Fares et al. 36 estimate an average price of 69 \$/kW based on data from the Federal Energy Regulatory Commission’s Form 1, where utilities can report their capital, operation, and maintenance costs related to grid reinforcement. In estimating this price, Fares et al. delineate distribution grid reinforcements from generation, transmission, and administrative costs. Rauschkolb et al. 37 use the same Federal Energy Regulatory Commission dataset to develop a linear model that accounts for both one-time capital costs of expanding grid capacity, and for increases in the annual costs of operating and maintaining a larger grid.
@@
Here, we modify the approach of Rauschkolb et al. to estimate a representative price of distribution grid reinforcement. We assume that grid capacities increase linearly from the business-as-usual capacities in 2025 to the all-electric capacities in 2050. We map future costs into net present costs using discounted cash flow analysis. Similar to Rauschkolb et al.’s approach 37, our modified approach accounts for both one-time capital costs of reinforcing infrastructure, and increased annual costs from operating and maintaining a larger grid. The Experimental Procedures section contains the details of our modifications, which yield a mean grid reinforcement price of 960 \$/kW and a 95% confidence interval of 587 to 1,331 \$/kW.
@@
The upper row of heat maps in Fig. 4 shows the estimated total distribution grid reinforcement cost for each of the lower 48 United States under low (left), medium (middle), and high (right) price scenarios. We compute these estimates by multiplying the grid reinforcement requirement, aggregated over all counties in the state, by prices of 587, 960, and 1,331 \$/kW, corresponding to the 2.5th, 50th, and 97.5th price percentiles. Unlike the per-household grid reinforcement requirement estimates in Fig. 3, the aggregate cost estimates in Fig. 4 reflect population density as well as load growth intensity. States with both high populations and high per-household grid reinforcement requirements have the highest total costs. For example, the 7.6 million households in New York, a densely populated state with cold winters, see a total cost of about \$56 billion, or about \$7,400 per household. Per-household grid reinforcement costs, shown in the bottom row of heat maps in Fig. 4, reach a maximum of \$19,000 per household under the 97.5th price percentile in the cold, sparsely populated state of North Dakota.
@@
Aggregated over the lower 48 United States, distribution grid reinforcement to accommodate 100% electrification of homes and private vehicles could cost \$320 to \$720 billion, or \$2,600 to \$5,800 per household. For context, residential energy bills cost Americans about \$230 billion (\$1,800 per household) per year 39. The Inflation Reduction Act of 2022 allocated about \$370 billion (\$2,900 per household) over ten years to spending on energy and climate 40. The United States annual military budget was about \$820 billion (\$6,500 per household) in 2023 41.
@@
Figure 5: Switching from resistance to heat-pump water heaters (top left) has little influence on grid reinforcement requirements. Decreasing thermal demand from buildings (top right) and switching from air- to ground-source heat pumps (bottom right) reduce grid reinforcement requirements significantly, particularly in cold climates. Coordinating device operation (bottom left) reduces grid reinforcement requirements in all climates.
The estimated distribution grid reinforcement requirements and costs in the previous sections come from an electrification scenario that extrapolates forward today’s housing stock, vehicle fleet, equipment configurations and efficiencies, and user behavior. Unmodified, these electrification choices could require \$320 to \$720 billion in distribution grid reinforcement costs nationwide. However, electrification with strategic demand-side management – referred to here as ‘smart electrification’ – could reduce these costs substantially.
To assess the cost reduction potential of smart electrification, we simulate four demand-side management strategies. First, the resistance water heaters simulated above could shift to heat-pump water heaters, which use about half the electricity per unit of heat output. Second, thermal demand from buildings could decrease due to better insulation and air sealing, due to shifts from detached single-family housing to less energy-intensive attached multifamily housing, or due to shifts to smaller housing. Third, software could reduce peak aggregate demand by coordinating the operation of water heaters, electric vehicles, and heat pumps. For example, heat pumps could preheat homes in anticipation of aggregate demand peaks, or electric vehicles could interleave their charging with heat pump operation. (We simulate device coordination through convex optimization with perfect information in one representative location per state, as discussed in the Supplementary Information document.) Fourth, the air-source heat pumps simulated above could shift to ground-source heat pumps, which use about one-third less electricity per unit of heat output and require less backup resistance heat.

### v3

@@
Replacing fossil-fueled appliances and vehicles with electric alternatives can reduce greenhouse gas emissions and air pollution in many settings. However, residential electrification can also raise electricity demand beyond the safe limits of electrical infrastructure. This can increase the risk of blackouts or require grid reinforcement that is often slow and expensive. Here, we estimate the physical and economic impacts on distribution grids of electrifying all housing and personal vehicles in each county of the lower 48 United States. We find that space heating is the main driver of grid impacts, with the coldest regions seeing demand peaks up to five times higher than today’s peaks. Accommodating electrification of all housing and personal vehicles is estimated to require 600 GW of distribution grid reinforcement nationally, at a cost of \$350 to \$790 billion, or \$2,800 to \$6,400 per household (95% confidence intervals). However, demand-side management could eliminate three-quarters of grid reinforcement costs.
@@
Distribution grids, Electrification, Heat pumps, Electric vehicles
@@ INTRODUCTION
Fossil-fueled driving, space heating, and water heating account for about one-third of United States greenhouse gas emissions 1. Replacing fossil-fueled vehicles and heaters with electric alternatives can significantly reduce these emissions today 2, and could essentially eliminate them if the United States achieves its goal of 100% carbon-free electricity by 2035 3. For these reasons, the Inflation Reduction Act of 2022 incentivizes adoption of electric vehicles and heat pumps (efficient electric heating and cooling machines) through tax credits and point-of-sale rebates 4, 5, 6. State and local governments have enacted complementary incentives, including some bans on fossil-fueled vehicles or heaters 7. Due in part to these policies, adoption of electric vehicles and heat pumps in the United States is accelerating 8.
@@
While electrification of fossil-fueled activities can deeply reduce emissions, it could also have serious impacts on the power grid. Today, electricity demand in most United States power systems peaks in the hottest hours of the year, driven mainly by air conditioning 9. Electrifying space heating can heighten annual demand peaks and shift them from summer to winter 10, 11, 12, 13, 14. Electric vehicle adoption is likely to further exacerbate demand peaks 15, 16, 17. Heightened demand peaks from electrification could cause power quality issues and make power outages more likely 18, 15. Mitigating these risks often requires replacing power lines and transformers, which are typically sized to accommodate power flows in near-worst-case demand forecast scenarios. Transformers in particular are already in short supply, with United States utilities experiencing lead times of up to two years and prices that have increased by 400 to 900% in the last three years 19. Utilities pass grid reinforcement costs on to ratepayers in the form of increased electricity prices 20, 21, which weaken economic incentives for further electrification.
@@
Past research on the grid impacts of electrification focused mainly on high-voltage transmission systems. At the nationwide scale, the National Renewable Energy Laboratory’s Electrification Futures Study 22, 10, 23 investigated how electrification might alter load profiles and transmission needs. In their high-electrification scenario (about 90% adoption of electric vehicles and 70% of heat pumps), peak electricity demand rose by 38%. A nationwide study from the Electric Power Research Institute 24 estimated that 75% adoption of electric vehicles and 50% of heat pumps (by floor area) would raise peak demand by 24% to 52%. Wilson et al. 2 used the ResStock tool 25 to investigate nationwide cost and emission impacts under various heat pump adoption scenarios. While Wilson et al. 2 did not explicitly consider grid impacts, they stressed the need for future work to understand how evolving load profiles should shape infrastructure investments. At the regional scale, an American Council for an Energy-Efficient Economy 26 study found that 30% adoption of EVs and 100% of heat pumps could increase peak demand in New England by 235%. White et al. 27 estimated that electrifying 100% of space heating would require 25% expansion of Texas grid capacity. Zhang et al. 28 estimated that 10% adoption of electric vehicles with uncoordinated charging would increase Midwest peak demand by 10%.
@@
In parallel with high-voltage transmission grid research at the nationwide and regional scales, several studies investigated medium-voltage distribution grid impacts in California specifically. Elmallah et al. 29 used the Distribution Deferral Opportunity Report, which the California Public Utilities Commission requires utilities to file each year, to analyze electrification impacts within the Pacific Gas & Electric service territory. Elmallah et al. estimated that distribution grid reinforcement to accommodate new load from electrification, including up to 100% adoption of electric vehicles and up to 50% adoption of heat pumps and water heaters, could cost \$1 to \$10 billion (about \$200 to \$2,000 per customer). Li et al. 30 used the same dataset to study 100% adoption of electric vehicles (but not heat pumps or water heaters) in California, estimating that 67% of distribution feeders could need reinforcement, at a total cost of \$6 to \$20 billion (about \$300 to \$900 per customer). Several other state-specific studies investigated the impacts of electrification on greenhouse gas emissions without considering impacts on power grids 14, 31, 32.
@@
As a complement to nationwide and regional transmission-level studies and California-specific distribution-level studies, this paper investigates the impacts of electrification on distribution grids nationwide. In addition to expanding the scope of distribution-level studies from the 58 counties in California to 3,100 counties in the lower 48 United States, this paper investigates strategies to mitigate grid impacts through demand-side management of space heating, water heating, and personal vehicles. These are the three main drivers of grid impacts from residential electrification, and we are aware of no distribution-level study that has considered demand-side management strategies for all of them. We find that almost all United States counties could require grid reinforcement in a 100% electrification scenario. On average over the three coldest climate zones (zones 5, 6, and 7), future peak demand is three times higher than today’s peak. Distribution grid reinforcement to accommodate 100% electrification nationwide could cost \$350 to \$790 billion, or about \$2,800 to \$6,400 per household. However, the demand-side management strategies evaluated here, such as improving equipment efficiencies and coordinating device operation, could reduce nationwide grid reinforcement costs by up to 71%.
We estimate distribution grid reinforcement requirements through bottom-up simulation. We estimate today’s distribution grid capacity in each county by simulating all the appliances and vehicles in 1,000 representative housing units during peak heating and cooling weeks, rescaling the overall peak aggregate demand to reflect the true number of housing units, and adding a safety margin to represent headroom in today’s capacity. We repeat the same process for an all-electric future. The estimated grid reinforcement requirement is the difference between the all-electric future grid capacity and today’s grid capacity. To simulate buildings and vehicles, we use the open-source EDGIE (Emulating the Distribution Grid Impacts of Electrification) toolbox 33. As illustrated in Fig. 1, EDGIE models heat pumps, electric vehicles, water heaters, and building thermal dynamics. It uses real data on weather 34, indoor temperature setpoints 35, domestic hot water use 36, and miscellaneous electrical loads 37. EDGIE implements empirically validated, multi-physics modeling at hourly or sub-hourly time resolution and county-level spatial resolution. We tune building, appliance, and vehicle parameters to county-specific data. We make all data publicly available to facilitate reproduction or extension of the work.
@@ RESULTS
Figure 2: In a hot week in Miami (left) and a cold week in Minneapolis (right), electricity demand peaks from 1,000 households exceed today’s peaks factors of two and four, respectively.
@@
We assess the physical impacts of home and vehicle electrification on distribution grids through bottom-up modeling. In each county, we simulate all the residential appliances and personal vehicles from 1,000 representative households, then rescale the aggregate load profile to reflect the true number of households. Simulated appliance and vehicle properties, housing sizes and types, and levels of insulation and air sealing all vary both across and within counties. We fit physical and behavioral parameters associated with buildings, thermal equipment, and vehicles to county- and climate-specific data, using real data wherever possible. The Experimental Procedures section and the Supplementary Information document provide more details on models and data sources.
We simulate peak heating and cooling weeks under two scenarios: (1) a business-as-usual scenario that reflects today’s building stock, appliances, and vehicles; and (2) an all-electric scenario that reflects a possible future with complete electrification of personal vehicles and residential space and water heating. In the all-electric scenario, each home has a central air-to-air heat pump with backup resistance heat. We compute design heating and cooling loads for each home under the 99% heating and 1% cooling design temperatures from the American Society of Heating, Refrigeration, and Air-Conditioning Engineers 2017 Handbook of Fundamentals and the Air Conditioning Contractors of America Manual J Design Conditions 8th Edition. Each central heat pump is sized to meet the larger of the home’s design heating load and design cooling load, up to a cap based on today’s equipment sizes. We model the largest available central heat pump as having five tons (17.6 kW) of nameplate cooling capacity. If the design heating or cooling load exceeds the maximum capacity of the largest available central heat pump under design conditions, we add a mini-split heat pump with one to three tons (3.5 to 10.6 kW) of nameplate cooling capacity. Resistance backup meets any heat demand in excess of the combined capacities of the central heat pump and the mini-split heat pump. In the two northernmost climate zones, we simulate adoption of heat pumps whose coefficients of performance meet the United States Department of Energy’s Cold Climate Heat Pump Challenge specifications 38. The modeled heat pump capacities and coefficients of performance vary with outdoor temperatures.
@@
Figure 3: Today (top left), per-household demand peaks are higher in hotter areas and in cold rural areas with significant amounts of electric heating. In an all-electric future (top right), per-household peaks increase and shift to colder areas. Per-household grid reinforcement requirements (bottom; the difference between the future peak and today’s peak, adjusted for distribution grid headroom) are highest in the Rocky Mountains, the upper Midwest, and the Northeast. Each county’s estimates are computed by simulating all devices in 1,000 representative households, then rescaling to the true number of households.
@@
The modeling approach outlined above enables analysis of the sizes and causes of likely grid impacts from electrification. Fig. 2 shows the aggregate power used by 1,000 households over a peak cooling week in Miami, Florida (left plot; a hot, humid climate) and a peak heating week in Minneapolis, Minnesota (right; cold). In these plots, the shaded areas from bottom to top represent miscellaneous electrical loads (blue; lights, computers, dishwashers, etc.), water heaters (orange), electric vehicles (yellow), and air-source heat pumps with backup resistance heat (purple). Dashed red lines show today’s annual demand peaks. In Miami, electrification doubles peak demand relative to today’s peak, due mainly to electric vehicle charging. In Minneapolis, peak demand increases by a factor of five, driven mainly by space heating in the coldest weather, when heat demand rises and heating efficiencies fall.
@@
We analyze grid impacts nationwide by running simulations for each county in the lower 48 United States. Fig. 3 shows heat maps of the estimated peak demand, normalized by the number of homes, in each county under the business-as-usual scenario (top left) and the all-electric scenario (top right). The top left plot shows that normalized peak demand (in units of kW per household) in the business-as-usual scenario tends to be higher in areas with more cooling demand, reaching a maximum of 13 kW per household in the hot, sunny Southwest. By contrast, the top right plot shows that normalized peak demand in the all-electric scenario is higher in colder areas, reaching a maximum of 25 kW per household in the northern Midwest.
We estimate the distribution grid reinforcement requirement as the difference between the all-electric and business-as-usual distribution grid capacities. We estimate the distribution grid capacities in each county and scenario by simulating the appliances and vehicles for 1,000 households, rescaling the aggregate demand to the true number of households, taking the 99th percentile, then adding a safety margin to reflect distribution grid headroom. We base headroom estimates on a hosting capacity analysis for seven utilities in New York State 39, which found that today’s headroom varies by utility from 15% to 36%. We use this headroom range for the business-as-usual scenario, randomizing over counties. For the all-electric scenario, we assume all counties have 20% headroom. The discussion section analyzes the sensitivity of grid reinforcement requirements to headroom assumptions in the business-as-usual and all-electric scenarios.
@@
The lower heat map in Fig. 3 shows the estimated grid reinforcement requirement, normalized by the number of households, required to accommodate 100% electrification in each county. The northern Midwest, Northeast, and Rocky Mountain regions require more grid reinforcement due to higher space heating demand. Grid reinforcement requirements reach a maximum of 23 kW per household in the northern Midwest. Many southern areas with mild winters require little or no grid reinforcement. In some cases, the lower heat map in Fig. 3 shows significant variation in grid reinforcement requirements between adjacent counties. This variation can be explained by differences in housing types and sizes, levels of insulation and air sealing, driving patterns, weather extremes, or other location-specific data. These geographic differences suggest a need for tailored infrastructure planning, as approaches that work for one county may not necessarily work for neighboring counties.
@@
To estimate the economic costs of reinforcing distribution grids to accommodate electrification, we multiply the aggregate grid reinforcement requirement for each county (in units of kW) by a \$/kW price. In reality, grid reinforcement prices vary significantly across the United States due to variation in grid topology and capacity, prevalence of overhead vs. underground power lines, costs of labor and equipment, and other factors. Due to a lack of reliable location-specific data on grid reinforcement prices, however, we use the same price distribution for each county.
@@
Figure 4: Statewide distribution grid reinforcement costs (top row) reflect both population densities and per-household grid reinforcement requirements. Per-household costs (bottom row) do not reflect population densities. Left column: Low-price scenario for grid reinforcement. Middle: Central price. Right: High price.
@@
We draw on three studies 29, 40, 41 to determine a plausible price range for typical distribution grid reinforcement, which may include installing power lines, transformers, switchgear, capacity banks, or other infrastructure. Based on Pacific Gas & Electric data, Elmallah et al. 29 estimated prices of 860 to 36,500 \$/kW, with the highest prices for the smallest projects. (We inflation-adjust all price data to 2024 dollars.) Zhang et al. 40 used Energy Information Administration data 42 to derive state-wise prices and a national average of 1,664 \$/kW, including a variety of administrative costs. Rauschkolb et al. 41 analyzed data from the Federal Energy Regulatory Commission’s Form 1 43, where utilities can report their capital, operation, and maintenance costs related to grid reinforcement. By contrast to prior studies, Raushkolb et al. separated the regular costs of sustaining an existing distribution grid from the incremental costs of expanding grid capacity to accommodate load growth. Rauschkolb et al. developed a linear model that accounts for both one-time capital costs from expanding grid capacity, and for increased annual costs from operating and maintaining a larger grid.
@@
Here, we modify Rauschkolb et al.’s approach to estimate a representative price range for distribution grid reinforcement. We assume grid capacities increase linearly from the business-as-usual capacities in 2025 to the all-electric capacities in 2050. We map future costs into net present costs using discounted cash flow analysis. Similar to Rauschkolb et al.’s approach, our modified approach accounts for both one-time capital costs of reinforcing infrastructure, and increased annual costs from operating and maintaining a larger grid. The Experimental Procedures section contains the details of our modifications, which yield a mean grid reinforcement price of 960 \$/kW and a 95% confidence interval of 587 to 1,331 \$/kW.
@@
The upper row of heat maps in Fig. 4 shows the estimated total distribution grid reinforcement cost for each of the lower 48 United States under low (left), central (middle), and high (right) price scenarios. We compute these estimates by multiplying the grid reinforcement requirement, aggregated over all counties in the state, by prices of 587, 960, and 1,331 \$/kW, corresponding to the 2.5th, 50th, and 97.5th price percentiles. Unlike the per-household grid reinforcement requirement estimates in Fig. 3, the aggregate cost estimates in Fig. 4 reflect population density as well as load growth intensity. States with both high populations and high per-household grid reinforcement requirements have the highest total costs. For example, the 7.6 million households in New York, a densely populated state with cold winters, see a total cost of about \$60 billion, or about \$7,900 per household. Per-household grid reinforcement costs, shown in the bottom row of heat maps in Fig. 4, reach a maximum of \$18,000 per household under the 97.5th price percentile in the cold, sparsely populated state of North Dakota.
The grid reinforcement cost estimates described above represent incremental spending, above and beyond the business-as-usual spending required to sustain existing distribution grids. The nationwide grid reinforcement cost estimate over 25 years in the central price scenario is \$569 billion. For comparison, we analyzed distribution grid spending data from the Federal Energy Regulatory Commission’s Form 1 43 from 1995 to 2019. Over that 25-year period, the total inflation-adjusted spending on distribution grid infrastructure – including all capital, operation, and maintenance costs associated with both sustaining existing grids and expanding capacity to accommodate load growth from residential, commercial, and industrial customers – was \$975 billion. Relative to this cost, the estimated incremental grid reinforcement cost of \$569 billion to accommodate home and vehicle electrification is a 58% increase. While this is not an exact comparison, it gives an approximate sense of scale for the incremental distribution grid spending required to accommodate residential electrification.
@@
Aggregated over the lower 48 United States, distribution grid reinforcement to accommodate 100% electrification of homes and private vehicles over 25 years is estimated to cost an additional \$350 to \$790 billion (\$110 to \$250 per household per year), above and beyond business-as-usual spending. For context, United States utilities spent \$975 billion on distribution grid infrastructure in the 25 years from 1995 to 2019 (\$390 per customer per year) 43. Residential energy bills cost Americans about \$230 billion annually (\$1,800 per household per year) 44. The Inflation Reduction Act of 2022 allocated about \$370 billion over ten years (\$290 per household per year) to spending on energy and climate 45. The United States annual military budget was about \$820 billion (\$6,500 per household per year) in 2023 46.
@@
Figure 5: Decreasing thermal demand from buildings (left) and switching from air- to ground-source heat pumps (right) reduce grid reinforcement requirements, particularly in cold climates. Coordinating device operation (middle) reduces grid reinforcement requirements in all climates.
The estimated distribution grid reinforcement requirements and costs in the previous sections come from an electrification scenario that extrapolates forward today’s housing stock, vehicle fleet, equipment configurations and efficiencies, and user behavior. Unmodified, these electrification choice are estimated to require \$350 to \$790 billion in distribution grid reinforcement costs nationwide. However, electrification with strategic demand-side management – referred to here as ‘smart electrification’ – could reduce these costs substantially.
To assess the cost reduction potential of smart electrification, we simulate three demand-side management strategies. First, thermal demand from buildings could decrease due to better insulation and air sealing, due to shifts from detached housing to less energy-intensive attached housing, and/or due to shifts to smaller housing. We simulate thermal demand reduction by increasing the effective thermal resistance (which models the combined effects of wall and roof insulation, window quality, outdoor air infiltration rates, and the surface area exposed to the outdoor air) by 25%, consistent with the ‘basic enclosure package’ from Maxim and Grubert 47. Second, the air-source heat pumps simulated above could shift to ground-source heat pumps, which use about one-third less electricity per unit of heat output and require less backup resistance heat. Third, software could reduce peak aggregate demand by coordinating the operation of water heaters, electric vehicles, and heat pumps. For example, heat pumps could preheat homes in anticipation of aggregate demand peaks, or electric vehicles could interleave their charging with heat pump operation. (We simulate device coordination through convex optimization with perfect information in one representative location per state, as discussed in the Supplementary Information document.) The demand-side management strategies simulated here are far from exhaustive. We leave for future work a range of measures, such as solar photovoltaics, bidirectional electric vehicle charging, home batteries, and thermal storage, that could further reduce grid reinforcement requirements.